# Simulating the economic impact of rationality through reinforcement learning and agent-based modelling

### v1

@@ 4 Results
Figure 2: Different emerging strategies for different economic environments. The middle panel (B) shows the time series of observed sales ($Y^{s}$) and price-deltas ($\tilde{\Delta}^{P}$) for a sample bounded rational agent (dashed line) and RL agent (solid line) under 3 different combinations of market competition ($z_{c}$) and degree of rationality ($N$) as indicated by the arrows. The 3 combinations give rise to 3 different strategies for the RL agents. From top to bottom, we find ‘perfect competition’, ‘dumping’ and ‘market power’ strategies (see main text for more details). The change in the emerging strategy for different economic environments is highlighted in the left (A) and right (C) panels, which plot the average value of price-delta (left axes) and sales (right axes) as a function of $z_{c}$ and $N$ respectively.
4.1 The microeconomic impact of rational agents
@@
We first evaluate the microeconomic impact of RL agents, which learn to efficiently produce and price goods in competition with each other and with bounded rational firms. In particular, for each firm we measure its sales $Y^{s}$ of consumption goods and its price delta $\tilde{\Delta}^{P}$, to assess the distance between the firm’s price and the average market price.
@@
In general, we find that RL agents choose and adapt their strategy according to the level of market competition $z_{c}$ and rationality $N$, outperforming the profits of bounded rational firms from the basis model.
Different emerging strategies for different levels of competition.
@@
We first evaluate the impact of different levels of competition through the model parameter $z_{c}$, considering a simulation with a single RL agent ($N=1$). Figure 2.A shows the agent policy in terms of price delta and sales. We observe that at increasing levels of market competition, the RL agent exhibits different strategies.
@@
When the competition in the market gets higher, for $z_{c}$ 3 or higher, the market power strategy becomes less profitable, and the learned price-quantity strategy changes abruptly. In these circumstances, the RL agent learns to drop the retail price below market level in order to undercut the competition, and it further learns that by charging lower prices it can produce and sell higher quantities. This, in turn, allows the RL agent to secure a large portion of the market share and accumulate high profits. We call this strategy a “dumping” strategy.
@@
For both price-quantity strategies, the RL agent has also successfully learned to avoid bankruptcy. This can be best observed from the time series of sales for the dumping strategy, since the sales of the RL agents remain high for the entire time while the sales of the bounded rational firm sometimes drop to zero for a certain number of steps, indicating a sudden bankruptcy and a subsequent slow recover.
Emerging perfect competition for multiple fully rational agents with a shared policy.
In Figure 2.C we investigate the case of multiple RL agents, with competition level fixed at $z_{c}=5$. In this setting, we train $N$ RL agents using a shared policy, i.e., all agents share and update the same Q-matrix.
@@
In Figure 2.B we show an insight into the simulation with $N=5$, showing the strategy of a sample RL agent, which typically sells fewer goods than a bounded rational agent, at similar or lower prices. It is again interesting to notice the adaptive and sophisticated behaviour of the RL agent in pricing goods from the top inset of Figure 2.B.
Figure 3: Independent RL-agents spontaneously segregate into strategic groups increasing overall profits. The left panel (A) shows the total mean cumulative rewards (i.e., profits) for a varying number of agents, with $z_{c}=5$, and for RL agents with shared or independent policies. Agents with independent policies always achieve higher overall rewards as a result of spontaneous segregation into strategic groups. This phenomenon can be clearly observed in the middle and right panels (B) and (C), showing the value of agent-specific price delta and sales as a function of a growing number of rational agents, with $z_{c}=5$. A small noise was added to the $x$-axis to better resolve very nearby points. The different strategic groups are particularly easy to spot when plotting price delta against sales, as done in the inset of (A) for $N=20$.
Finally, we note that average profits of RL agents diminish substantially in the transition from a low rationality environment ($N=1,2$) which can be exploited with a dumping strategy, to a high rationality environment ($N\geq 3$) where RL agents are forced into perfect competition to stay profitable and bankruptcy-free. This abrupt change in profits is reported in Figure 3.A. The figure also illustrates that profits keep decreasing although more gradually also for $N\geq 3$, in line with classical microeconomic theory. Although we do not explicitly show the profits/rewards of bounded rational firms, we note here that these are consistently lower than those of RL agents as illustrated for a single run in the right panel of Figure 1.
@@
  • Market power. Firms maximise profits by exploiting the imperfection of the market, with households being able to visit only a small number of firms for their consumption. They produce and sell small amounts of goods but they charge a very high price, thus at the same time maximising revenues and minimising costs.
@@
Independent RL agents increase profits over perfect competition by segregation.
We now investigate the case of multiple RL agents trained with independent Q-matrices, and hence free to learn different policies. Figure 3.A shows the cumulative average rewards achieved by an increasing number $N$ of RL agents at fixed $z_{c}=5$, for training with both independent and shared policies. The first immediately visible emergent phenomenon is the fact that RL agents with individual policies achieve higher total profits/rewards on average as compared with RL agents that are constrained to the same common policy. While this may seem counterintuitive, since an agent would maximise its individual rewards at the expense of the others, we recall that our model is not a zero-sum game. Independent agents can achieve higher profits by adjusting their strategies and holding different market niches. Specifically, we here observe an interesting phenomenon of spontaneous segregation of the RL agents into different strategic groups, where agents within each group follow one of the three strategies described above. This, in turn, increases profits substantially over the shared policy scenario since certain strategic groups can still carry on exploitative behaviour which would not be sustainable for all RL agents at the same time. The spontaneous segregation of agents is particularly clear from the inset of Figure 3.A, where the 20 RL agents are seen to cluster into all three discovered strategies according to their median sales and price.
@@
Figure 4: RL agents always increase total macroeconomic output, but only perfect competition also increases economic stability. Mean (A), and standard deviation (B) of the GDP in the last 1000 steps of simulations with a growing number of rational agents, trained either with shared or independent policies. Error bars represent the standard error on average quantities. The last panel (C) shows the response function of consumption, GDP and GDP deflator to a positive shock in the propensity to consume of all households. The impulse responses are shown for the basis model and for the models with shared and independent policies with $N=7$. All results of the figure are obtained with $z_{c}=5$.
4.2 The macroeconomic impact of rational agents
@@
Increased rationality implies higher overall output.
Figure 4.A shows the average Gross domestic product (GDP) of simulations with RL agents with shared or independent policies for increasing $N$. The figure clearly shows that, independently of the type of policy, any level of increased rationality $N$ gives rise to a significant increase in the overall output of the economic system with respect to the basis model. This is not surprising since rational firms have a greater incentive to increase production levels to increase profits. This happens both in ‘exploitable’ economic environments that allow few RL agents to gain a high market share via dumping ($N=1,2$) and in ‘unexploitable’ markets where RL agents are forced into perfect competition and achieve higher outputs collectively.
Increased rationality implies higher economic stability only when it also implies perfect competition.
Figure 4.B shows the economic instability of the R-MABM, measured by the standard deviation of the GDP, for an increasing number $N$ of RL agents trained with shared or independent policies. While overall output always increases with increasing rationality, economic stability does not necessarily increase.
Specifically, we find that rationality increases economic stability only when it also implies the adoption of the perfect competition strategy by rational firms. On the contrary, rational firms adopting an exploitative behaviour, like the dumping strategy, capitalise on firms with limited rationality making them more susceptible to bankruptcy, thus increasing overall economic instability.
@@
Perfect competition gives rise to the highest overall output and to efficient responses to aggregate shocks.
The highest overall outputs are achieved at $N=7,10,15$ for RL agents with shared policy following a perfect competition strategy. This result is not trivial, as one might imagine the dumping strategy to lead to greater outputs as it implies the highest production levels for rational firms. The perfect competition strategy also appears to give rise to quick and effective responses to aggregate shocks, as exemplified in Figure 4.C. The figure shows the impulse response functions for three macroeconomic variables: consumption, real GDP, and GDP deflator (a measure of inflation). The impulse (or ‘shock’) consists of an instantaneous 30% increase in the propensity to consume of all households, and is computed for the baseline model and for R-MABM models with $N=7$ and shared/independent policies. The curves are obtained following the technique described in [15].
@@ 5 Conclusions
We introduced an innovative and promising application for multi-agent reinforcement learning: the extension of traditional macroeconomic agent-based models (ABMs) with ‘fully rational’ agents that optimise their behaviour to maximise a reward function. Specifically, in this work we apply this idea to a classical macroeconomic ABM, proposing a ‘Rational macro ABM’ (R-MABM) by substituting bounded rational firms in the classical model with profit-maximising reinforcement learning (RL) agents. Our R-MABM model allows the study of a variety of issues in economics that cannot be investigated with traditional macro ABMs, where agents simply follow fixed behavioural rules. We showcase how the R-MABM allows a thorough examination of the impact of increased levels of rationality (modelled as the number of RL agents) in the economy. We find a number of empirical results that can be well connected with ideas and concepts from economic theory.
@@
On the microeconomic level, we find that RL agents (i.e., fully rational firms) spontaneously learn three distinct price-quantity strategies to maximise profits depending on the economic environment they are immersed in. When competition is very low, they learn a ‘market power’ strategy consisting of charging very high prices and producing very little, relying on the fact that with low competition there is always demand for their goods. When competition is higher but there are few rational firms, they learn a ‘dumping’ strategy consisting of flooding the market with high quantities of goods sold at very low prices, thus undercutting competition from the bounded rational agents. Finally, when both competition and the number of rational firms in the economy are high, RL agents are forced into a ‘perfect competition’ strategy consisting of producing enough goods to satisfy the demand and selling them at market prices. Notably, we also find that when RL agents are allowed to follow independent policies they spontaneously learn to segregate into different strategic groups thus increasing profits for all.
@@
On the macroeconomic level, we find that rationality always improves overall economic output, but that macroeconomic stability is improved only when RL agents are forced into a strategy of perfect competition. We further find that RL agents in perfect competition strategy typically give rise to the greatest economic output and to high economic stability and resilience to shocks.
@@
This work represents an important first step in the investigation of the use of RL in traditional economic ABMs and, more generally, an important step towards the integration of artificial intelligence in economic simulation models. Our R-MABM framework allows for stable multi-agent learning, and thus it provides a solid foundation for many possible extensions. Future work could involve increasing the action space of rational firms to include investment choices, or studying the effect of the introduction of RL agents in place of other agent classes in the basis model, such as banks and households. Finally, it would be interesting to add an RL government or central bank to study the emergence of optimal fiscal or monetary policies.
@@
In the interest of reproducibility, the code to simulate the R-MABM framework will be released in open source soon. In the meantime, please do not hesitate to get in touch if you are interested.

### v2

@@ 2.2. The RL agents
Figure 2. Different emerging strategies for different economic environments. The middle panel (B) shows the time series of observed sales ($Y^{s}$) and price-deltas ($\tilde{\Delta}^{P}$) for a sample bounded rational agent (dashed line) and RL agent (solid line) under 3 different combinations of market competition ($z_{c}$) and degree of rationality ($N$) as indicated by the arrows. The 3 combinations give rise to 3 different strategies for the RL agents. From top to bottom, we find ‘perfect competition’, ‘dumping’ and ‘market power’ strategies (see main text for more details). The change in the emerging strategy for different economic environments is highlighted in the left (A) and right (C) panels, which plot the average value of price-delta (left axes) and sales (right axes) as a function of $z_{c}$ and $N$ respectively.
3. Experimental setup
We simulate the macroeconomic model considering a single bank, $1000$ workers, $20$ K-firms, and $100$ C-firms. In our experiments, we vary the levels of competition ($z_{c}$) from $2$ to $10$, and the number of RL agents ($N$), representing fully rational C-firms, from $1$ to $20$. Since we focus exclusively on C-firms in our experiment, we will refer to them simply as ‘firms’ for the remainder of the work. All simulation parameters are described in Table 1, while additional parameters of the basis model are those of Assenza et al. (Assenza et al., 2015).
We train the RL agents for each model configuration, considering both shared and independent policies. We consider $T_{\textit{train}}=100$ episodes for the training phase, a number which we have found more than sufficient for reaching a full convergence of the Q-learners in all our experiments, and consider $T_{\textit{test}}=20$ episodes for testing. Unless otherwise specified, we report averages of the simulation results over the 20 testing episodes, with error bars representing one standard deviation.
Similarly to some existing MARL frameworks for economic modelling, we consider a sort of curriculum learning (Bengio et al., 2009): the simulation starts without the RL agents, which are introduced after $t_{\textit{burn-in}}$ steps. Then, the RL agents assume control of $N$ firms for additional $t_{\textit{sim}}$ steps. At the start of the training phase, RL agents are initialised with an empty Q-matrix i.e. all values are zero. During training, and at each episode $\tau=1,\dots,T_{\textit{train}}$ , the agents update the exploration rate $\epsilon$ of the $\epsilon$-greedy policy in Eq. (15) as $\epsilon\leftarrow\max\left(0.9^{\tau-1},0.01\right)$.
4. Results
We now discuss the simulations obtained with our R-MABM framework, with standard firms gradually substituted by RL agents trained to maximise profits. We divide our results into the microeconomic effects (Section 4.1) and the macroeconomic effects (Section 4.2) of the introduction of rational RL agents.
4.1. The microeconomic impact of RL agents
@@
We first evaluate the microeconomic impact of RL agents learning to efficiently produce and price goods in competition with each other and with bounded rational firms. For each firm, we measure its sales $Y^{s}$ of consumption goods and its price delta $\tilde{\Delta}^{P}$, to assess the distance between the firm’s price and the average market price.
@@
In general, we find that RL agents adapt their strategy to the level of market competition $z_{c}$ and rationality $N$, outperforming the profits of bounded rational firms from the basis model.
@@
Different emerging strategies for different levels of competition. We first evaluate the impact of different levels of competition through the model parameter $z_{c}$, considering a simulation with a single RL agent ($N=1$). Figure 2.A shows the agent policy in terms of price delta and sales. We observe that at increasing levels of market competition, the RL agent exhibits different strategies.
@@
When the competition in the market gets higher, for $z_{c}\geq 3$, the market power strategy becomes less profitable, and the learned price-quantity strategy changes abruptly. In these circumstances, the RL agent learns to drop the retail price below market level in order to undercut the competition, and it further learns that by charging lower prices it can produce and sell higher quantities. This, in turn, allows the RL agent to secure a large portion of the market share and accumulate high profits. We call this strategy a “dumping” strategy.
Figure 3. Independent RL-agents spontaneously segregate into strategic groups increasing overall profits. The left panel (A) shows the total mean cumulative rewards (i.e., profits) for a varying number of agents, with $z_{c}=5$, and for RL agents with shared or independent policies. Agents with independent policies always achieve higher overall rewards as a result of spontaneous segregation into strategic groups. Segregation can be clearly observed in the middle and right panels (B) and (C), showing the value of agent-specific price delta and sales as a function of a growing number of rational agents, with $z_{c}=5$. A small noise was added to the $x$-axis to better resolve very nearby points. The different strategic groups are particularly easy to spot when plotting price delta against sales, as done in the inset of (A) for $N=20$.
@@
For both strategies, the RL agent has also successfully learned to avoid bankruptcy. This can be best observed from the time series of sales for the dumping strategy, as the sales of the RL agents remain high for the entire time while the sales of the bounded rational firm sometimes drop to zero for a certain number of steps, indicating a sudden bankruptcy and a subsequent slow recover.
Emerging perfect competition for multiple fully rational agents with a shared policy. In Figure 2.C we investigate the case of multiple RL agents, with competition level fixed at $z_{c}=5$. In this setting, we train $N$ RL agents using a shared policy, i.e., all agents share and update the same Q-matrix.
@@
In Figure 2.B we show an insight into the simulation with $N=5$, showing the strategy of a sample RL agent, which typically sells fewer goods than a bounded rational agent, at similar or lower prices. It is again interesting to notice the adaptive pricing behaviour of the RL agent from the top inset of Figure 2.B.
Finally, we note that average profits of RL agents diminish substantially in the transition from an environment with $N=1,2$, which can be exploited with a dumping strategy, to an environment with $N\geq 3$, where RL agents are forced into perfect competition to stay profitable and bankruptcy-free. This abrupt change in profits is reported in Figure 3.A. The figure also illustrates that profits keep decreasing although more gradually also for $N\geq 3$, in line with classical microeconomic theory. Although we do not explicitly show the profits/rewards of bounded rational firms, we note here that these are consistently lower than those of RL agents as illustrated for a single run in the right panel of Figure 1.
@@
  • Market power. Firms maximise profits by exploiting the imperfection of the market, with households being able to visit only a small number of firms for their consumption. They produce and sell small amounts of goods charging very high prices, thus at the same time maximising revenues and minimising costs.
@@
Figure 4. RL agents always increase total macroeconomic output, but only perfect competition also increases economic stability. Mean (A), and standard deviation (B) of the GDP in the last 1000 steps of simulations with a growing number of rational agents, trained either with shared or independent policies. Error bars represent the standard error on average quantities. The last panel (C) shows the response function of consumption, GDP and GDP deflator to a positive shock in the propensity to consume of all households. All results of the figure are obtained with $z_{c}=5$ and the impulse responses are obtained with $N=7$.
Independent RL agents increase profits over perfect competition by segregation. We now investigate the case of multiple RL agents trained with independent Q-matrices, and hence free to learn different policies. Figure 3.A shows the cumulative average rewards achieved by an increasing number $N$ of RL agents at fixed $z_{c}=5$, for training with both independent and shared policies. The first immediately visible emergent phenomenon is the fact that RL agents with individual policies achieve higher total profits/rewards on average as compared with RL agents that are constrained to the same common policy. While this may seem counterintuitive, since an agent would maximise its individual rewards at the expense of the others, we recall that our model is not a zero-sum game. Independent agents can achieve higher profits by adjusting their strategies and holding different market niches. Specifically, we here observe an interesting phenomenon of spontaneous segregation of the RL agents into different strategic groups, where agents within each group follow one of the three strategies described above. This, in turn, increases profits substantially over the shared policy scenario since certain strategic groups can still carry on exploitative behaviour which would not be sustainable for all RL agents at the same time. The spontaneous segregation of agents is particularly clear from the inset of Figure 3.A, where the 20 RL agents are seen to cluster into all three discovered strategies according to their median sales and price.
@@
4.2. The macroeconomic impact of RL agents
@@
Increased rationality implies higher overall output. Figure 4.A shows the average gross domestic product (GDP) of simulations with RL agents with shared or independent policies for increasing $N$. The figure clearly shows that, independently of the type of policy, any number of RL agents $N$ gives rise to a significant increase in the overall output of the economic system with respect to the basis model. This is not surprising since rational firms have a greater incentive to increase production levels to increase profits. This happens both in ‘exploitable’ economic environments that allow few RL agents to gain a high market share via dumping ($N=1,2$) and in ‘unexploitable’ markets where RL agents are forced into perfect competition and achieve higher outputs collectively.
Increased rationality implies higher economic stability only when it also implies perfect competition. Figure 4.B shows the economic instability of the R-MABM, measured by the standard deviation of the GDP, for an increasing number $N$ of RL agents trained with shared or independent policies. While overall output always increases with increasing rationality, economic stability does not necessarily increase. Specifically, we find that rationality increases economic stability only when it also implies the adoption of the perfect competition strategy by RL firms. On the contrary, RL firms adopting an exploitative behaviour like the dumping strategy, capitalise on firms with limited rationality making them more susceptible to bankruptcy, thus increasing overall economic instability.
@@
Perfect competition gives rise to the highest overall output and to efficient responses to aggregate shocks. The highest overall outputs are achieved at $N=7,10,15$ for RL agents with shared policy following a perfect competition strategy, in spite of the fact that the dumping strategy implies higher production levels for rational firms. The perfect competition strategy also appears to give rise to quick and effective responses to aggregate shocks, as exemplified in Figure 4.C. The figure shows the impulse response functions for three macroeconomic variables: consumption, real GDP, and GDP deflator (a measure of inflation). The impulse (or ‘shock’) consists of an instantaneous 30% increase in the propensity to consume of all households and is computed for the baseline model and for R-MABM models with shared/independent policies. The curves are obtained following the technique described in (Gatti and Grazzini, 2020).
@@ 5. Conclusions
In this work, we introduce the use of multi-agent reinforcement learning to extend traditional macroeconomic agent-based models (ABMs). Specifically, we propose a ‘Rational macro ABM’ (R-MABM) framework that allows substituting firms in a classical macro ABM with profit-maximising reinforcement learning (RL) agents. We define RL agents to be ‘fully rational’ as they explicitly optimise a reward function, and contrast them with ‘bounded rational’ agents with fixed behavioural rules, hence the name of the framework. We showcase how our framework allows for modelling and studying increased levels of rationality (the number of RL agents) in the economy, a problem that cannot be investigated with traditional macro ABMs.
@@
We find a number of empirical results that can be well connected with ideas and concepts from economic theory. On the microeconomic level, we find that RL agents spontaneously learn three distinct price-quantity strategies to maximise profits depending on the economic environment they are immersed in. When competition is very low, they learn a ‘market power’ strategy consisting of charging very high prices and producing very little, relying on the fact that with low competition there is always demand for their goods. When competition is higher but there are few RL firms, they learn a ‘dumping’ strategy consisting of flooding the market with high quantities of goods sold at very low prices, thus undercutting competition from the bounded rational agents. Finally, when both competition and the number of RL firms in the economy are high, RL agents are forced into a ‘perfect competition’ strategy consisting of producing enough goods to satisfy the demand and selling them at market prices.
Notably, we also find that when RL agents are allowed to follow independent policies they spontaneously learn to segregate into different strategic groups thus increasing profits for all.
@@
On the macroeconomic level, we find that RL firms always improve overall economic output, but that macroeconomic stability is improved only when they are forced into a strategy of perfect competition. We further find that RL agents in perfect competition strategy typically give rise to the greatest economic output and to high economic stability and resilience to shocks.
@@
This work is a first step towards the investigation of the use of RL in traditional economic ABMs and, more generally, an important step towards the integration of artificial intelligence in economic simulation models. Our R-MABM framework allows for stable multi-agent learning, and thus it provides a solid foundation for many possible extensions. Future work could involve increasing the action space of rational firms to include other choices such as investments, or studying emergent decision-making from RL models of other agent classes in the basis model, such as the government, the central bank, the commercial banks or the households.
Acknowledgements.
We would like to thank Adrián Carro (Bank of Spain), Herbert Dawid (Bielefeld University) and Marco Pangallo (CENTAI institute) for constructive feedback on this work. The views and opinions expressed in this paper are those of the authors and do not necessarily reflect the official policy or position of Banca d’Italia.
@@ Acknowledgements.
In the interest of reproducibility, the code to simulate the R-MABM framework is available in open source at https://github.com/Brusa99/R-MABM. The code is based on the MABM model as available at https://github.com/bancaditalia/ABCredit.jl.